OpenStudy (anonymous):

how to solve this quadratic equation 10z2+19z+6=0

OpenStudy (anonymous):

you need to find factors of 10 and 6 which when added give 19 5 * 3 and 2 * 2 will fit the bill: so ( 5x + 2)(2x + 3) = 0 5x + 2 = 0 or 2x + 3 = 0 x = -2/5 or x = -3/2 solution set is -2/5 , -3/2

OpenStudy (anonymous):

thanks for replying sir.

OpenStudy (anonymous):

yw